A ball is thrown obliquely upward with the initial speed 12m / s and with the elevation angle 42 degrees. what speed does the ball have after 0.18s?
1
Expert's answer
2019-05-17T11:17:29-0400

Horizontal component is


"v_x=v \cos{42}"

Vertical component is


"v_y=v \sin{42}-gt"

Thus,


"v=\sqrt{(12\cos{42})^2+(12\sin{42}-(9.8)(0.18))^2}=11\frac{m}{s}"